OpenStudy (anonymous):

In the slope intercept form, write the equation of the line with slope -1/2 passing through (0,5). y=mx+b

OpenStudy (wwwhhhaaattt?):

So you do need more help.

OpenStudy (anonymous):

Yeah, I didn't see this paper lol

OpenStudy (wwwhhhaaattt?):

Lol Ok So what do you think it is.

OpenStudy (studygurl14):

slope-intercept form y= mx + b where m=slope

OpenStudy (studygurl14):

What's the first step?

OpenStudy (anonymous):

I'm not really sure

OpenStudy (wwwhhhaaattt?):

Damit StudyGurl WHY DO YOU HAVE TO BE SO SMART

OpenStudy (anonymous):

Lol

OpenStudy (studygurl14):

quote "line with slope -1/2"

OpenStudy (wwwhhhaaattt?):

Shes THE SMARTEST PERSON IVE EVER MEET.

OpenStudy (studygurl14):

lol @WWWHHHAAATTT? I'm not. I've just taken A LOT of math :)

OpenStudy (anonymous):

y-5=-1/2(x-0) when moved into slope intercept it means that y=-1/2x+5

OpenStudy (anonymous):

Okay

OpenStudy (studygurl14):

That's another way to do it @WinMath which totally works. I like that method :)

OpenStudy (wwwhhhaaattt?):

@StudyGurl14 I'm SSSSSSSSSSSSOOOOOOOOOOOOO Sorry

OpenStudy (studygurl14):

Sorry for what?

OpenStudy (anonymous):

I would appreciate a medal :P

OpenStudy (anonymous):

No problem

OpenStudy (studygurl14):

Hey @WWWHHHAAATTT? I'll give you a medal if you give me a medal...

OpenStudy (wwwhhhaaattt?):

There

OpenStudy (studygurl14):

thx